Calculate the average induced voltage between the tips of the wings of a Boeing 747 flying at 910 km/hr above East Lansing. The downward component of the earth's magnetic field at this place is 0.7·10-4 T. (DATA: Assume that the wingspan is 60 meters.)
Answer #1

Let us

From Faraday's law, induced emf

Emf=-N\frac{\Delta \Phi }{\Delta T}

where..

N=no of turns,

\Phi=BA=magnetic flux

B=external magnetic field

A=area of coil

Emf=-\Delta \left ( \frac{B\times T}{T} \right ).................where (n=1 span)


In this case the emf is due to a changing sweep area ∆A/T
910km/hr = 252.778m/s
\Delta \frac{A}{T}=252.778m/s\times 60min=15166.2m^{2}/s

Emf=B\times \Delta \frac{A}{T}

Emf=0.7\times10^{-4}T \times 15166.2m^{2}/s

Emf=1.061V

Average indused voltage is 1.061V